OpenStudy (anonymous):

Medal for whoever helps me first with all my questions. (: 13.Eric knew his trip to Lawrenceburg was 548 miles. If he drove for 8.9 hours, what mph did he average on his trip? You may use the formula d = rt. Round your answer to the nearest whole number. A.60 mph B.61 mph C.62 mph 14.Murray's stocks changed these amounts over these months: -8.23 in September, -5.67 in October, 4.8 in November, and -2 in December. What is the average change over that time period? A.- 2.775 B.+ 2.775 C.+ 5.175 D.- 5.175

OpenStudy (anonymous):

1.C

OpenStudy (anonymous):

13 b 14 b

OpenStudy (anonymous):

2.B

OpenStudy (anonymous):

13 is 61.57 so 62 tat is c 14 is b

OpenStudy (anonymous):

13. v=d/t=548/8.9=61.5=62 mph 14. [(-8.23)+(-5.67)+(4.8)+ (-2)]/4= -2.775

OpenStudy (anonymous):

nice medal for free answer